OpenStudy (hy123):

(9(sqrt)6+3(sqrt)3)(6(sqrt)6-8(sqrt)3) multiply

OpenStudy (hy123):

\[(9\sqrt{6}+3\sqrt{3})(6\sqrt{6}-8\sqrt{3})\]

OpenStudy (hy123):

is \[252-162\sqrt{2} ?\]

OpenStudy (anonymous):

\[9\sqrt{6}(6\sqrt{6}-8\sqrt{3})+3\sqrt{3}(6\sqrt{6}-8\sqrt{3})\]

OpenStudy (anonymous):

54*6-72*sqrt(18)+18*sqrt(18)-24*3 324-54*sqrt(18)-72 252-(162*sqrt(2))

OpenStudy (anonymous):

ya right your answer

OpenStudy (anonymous):

@hy123